# Conformal prediction for causal effects of continuous treatments

### v3

@@ 2 Related Work
Several extensions have been developed for CP. One literature stream focuses on CP with marginal coverage under distribution shifts between training and test data [e.g., 8, 11, 15, 16, 17, 18, 19, 36, 42, 50, 57]. Our setting later also involves a distribution shift due to the intervention on the treatment, but differs from the latter in that the true distribution shift is unknown. Another literature stream constructs intervals conditional on the variables following the shifted distribution. Since, in general, exact conditional coverage has been proven impossible [35, 52], the works in this literature stream have two key limitations: (1) they only guarantee approximate conditional coverage [e.g., 5, 7, 35, 44]; or (2) they are restricted to specific data structures such as binary variables [e.g., 35, 52]. Because of that, none of the existing methods for marginal and conditional coverage can be applied to derive prediction intervals with finite-sample validity guarantees for causal quantities of continuous treatments.
@@
Conformal prediction for causal quantities: Only a few works focus on CP for causal quantities (see Fig. 2). Examples are methods aimed at off-policy learning [49, 60], conformal sensitivity analysis [58], or meta-learners for the conditional average treatment effect (CATE) [2, 27, 36, 54]. However, there are crucial differences to our setting: First, the existing works (a) assume that the propensity is known and thus achieve finite-sample coverage guarantees, or the existing works (b) focus on the easier task of giving asymptotic guarantees but then might suffer from under-coverage because of which the intervals are not faithful. Only Lei and Candès [36] provides finite-sample coverage guarantees under estimated propensity scores. However, all existing CP methods are designed for binary or discrete treatments. Applying such methods to discretized continuous treatments leads to ill-defined causal estimands. Therefore, none of the existing methods are applicable to our continuous treatment setting. We offer a detailed discussion in Supplement E.
@@ 3 Problem formulation
Second, the propensity score $\pi$ is commonly unknown in observational data and, therefore, must be estimated, which introduces additional uncertainty that one must account for when constructing CP intervals ($\rightarrow$ Challenge  b). Crucially, existing coverage guarantees [e.g., 52, 50] do not hold in our setting. Instead, we must derive new intervals with valid coverage under propensity shift.
@@ 4 CP intervals for potential outcomes of continuous treatments
$\hat{\theta}_{S}:=\arg\min_{\theta\in\mathbb{R}}\frac{1}{n-m}\left(\sum_{i=m+1}^{n}l_{\alpha}(\theta,S_{i})+l_{\alpha}(\theta,S)\right),$ (3)
@@
$l_{\alpha}(\theta,S):=\begin{cases}(1-\alpha)(S-\theta),&\text{ if }S\geq\theta,\\ \alpha(\theta-S),&\text{ if }S<\theta.\end{cases}$ (4)
@@ 4.1 Scenario 1: Known propensity score
Let $\mathcal{F}$ define a finite-dimensional function class that includes the function $f$ characterizing the shift in the (potentially unknown) propensity function $\pi$ (see Eq. (2)). Define the distribution-shift-calibrated $(1-\alpha)$-quantile of the non-conformity scores as
$\hat{g}_{S}(X_{n+1}):=\arg\min\limits_{g\in\mathcal{F}}\frac{1}{n-m}\left(\sum_{i=m+1}^{n}l_{\alpha}(g(X_{i}),S_{i})+l_{\alpha}(g(X_{n+1}),S)\right)$ (5)
@@
Building upon Lemma 4.1, we derive our first main result in Thm. 4.2. We define the finite-dimensional function class of interest as :=$\mathcal{F}:=\{\theta\frac{\pi(a+\Delta_{A}\mid x)}{\pi(a\mid x)}\mid\theta\in\mathbb{R}^{+}\}$. It is easy to verify that all $f\in\mathcal{F}$ represent the desired propensity shift to $\tilde{\pi}(a\mid x)=\pi(a+\Delta_{A}\mid x)$ as defined in Eq. 2.
@@
Consider a new datapoint with $X_{n+1}=x_{n+1}$, $A_{n+1}=a_{n+1}$, and $A^{\ast}(X_{n+1})=a^{*}=a_{n+1}+\Delta_{A}$. Let $\eta^{S}=\{\eta_{m+1}^{S},\ldots,\eta_{n+1}^{S}\}\in\mathbb{R}^{n+1}$ be the optimal solution to
@@
[…]
@@
We provide a full proof in Supplement D.2. Here, we briefly outline the underlying idea of the proof. First, we show that the function class $\mathcal{F}$ indeed satisfies Eq. (2) for the intervention $A^{*}(X)=A+\Delta_{A}$, and we then rewrite Eq. (5) as a convex optimization problem. Next, we exploit the strong duality property. We optimize over the corresponding dual problem to receive a dual prediction set with equal coverage probability. Finally, we derive $S^{\ast}$ from the dual prediction set to construct $C_{n+1}$ and prove the overall coverage guarantee. ∎
@@ 4.2 Scenario 2: Unknown treatment policy
$\tilde{\pi}(a\mid x)=\frac{\delta_{a^{\ast}}(a)}{\hat{\pi}(a\mid x)}\frac{\hat{\pi}(a\mid x)}{\pi(a\mid x)}\pi(a\mid x)=c_{a}\frac{\delta_{a^{\ast}}(a)}{\hat{\pi}(a\mid x)}\pi(a\mid x)=\frac{f(a,x)}{\mathbb{E}_{P}[f(A,X)]}\pi({a\mid x}),$ (13)
@@
:=$\mathcal{F}:=\Bigg{\{}\frac{c_{a}}{\sqrt{2\pi}\sigma}\frac{\exp(-\frac{(a-a^{\ast})^{2}}{2\sigma^{2}})}{\hat{\pi}(a\mid x)}\Bigm{|}0<\sigma,\frac{1}{M}\leq c_{a}\leq M\Bigg{\}}.$ (16)
@@
can be represented in terms of Eq. (2) through functions $f\in\mathcal{F}$.
@@
$\min\limits_{{\sigma>0,\frac{1}{M}\leq c_{a}\leq M}}\quad\sum_{i=m+1}^{n+1}(1-\alpha)u_{i}+\alpha v_{i}$ (PS)
@@
for the imputed score $S_{n+1}=S$. As the score is unknown, computing the CP interval would require solving (PS) for all $S\in\mathbb{R}$, yet which is computationally infeasible. As a remedy, we previously exploited properties of the dual optimization problem and the Lagrange multipliers of the convex problem in Theorem 4.2 to efficiently compute the CP intervals (see the proof of Theorem 4.2 in Supplement D). However, the present non-convex problem does not automatically allow for the same simplifications. Instead, we now present a remedy for efficient computation of the CP intervals in the following lemma. We prove Lemma 4.4 in Supplement D.
@@
The problem (PS) is Type-I invex and satisfies the linear independence constraint qualification (LICQ).
Lemma 4.4 allows us to derive properties of the present non-convex optimization problem in terms of the Karush-Kuhn-Tucker (KKT) conditions. For this, we note that the fulfillment of the LICQ serves as a sufficient regularity condition for the KKT to hold at any (local) optimum of (PS). Combined with the Type-I invexity of the objective function and the constraints, the KKT conditions are not only necessary but also sufficient for a global optimum. As a result, we can employ the KKT conditions at the optimal values [^6] We provide an interpretation of the optimal values in Supplement F. Therein, we further discuss the implications of the proposed kernel smoothing of $\delta_{a^{\ast}}(a)$. $\sigma^{\ast}$ and $c_{a}^{\ast}$ to derive coverage guarantees of our CP interval in a similar fashion as in Theorem 4.2. We thus arrive at the following theorem to provide CP intervals for the scenario with unknown propensity scores.
@@ 5 Experiments
Baselines: As we have discussed above, there are no baselines that directly compute prediction intervals with finite-sample guarantees for potential outcomes of continuous treatments. Therefore, we compare our method against MC dropout [14] and deep ensemble methods [33]. Yet, we again emphasize that MC dropout is an ad hoc method with poor approximations of the posterior, which is known to give unfaithful intervals [34]. Furthermore, we report the empirical coverage achieved by intervals from a Gaussian process regression (GP). By doing so, we consider a method that assesses the underlying aleatoric uncertainty.
Implementation: All methods are implemented with $\phi$ as a multi-layer perceptron (MLP) and an MC dropout regularization of rate 0.1. Crucially, we use the identical MLP for both our CP method and MC dropout. Hence, all performance gains must be attributed to the coverage guarantees of our conformal method. In the MC dropout baseline, the uncertainty intervals are computed via Monte Carlo sampling. In scenario 2, we perform conditional density estimation by conditional normalizing flows [51]. Implementation and training details are in Supplement F.
Performance metrics: We evaluate the methods in terms of whether the prediction intervals are faithful [e.g., 21]. That is, we compute whether the empirical coverage of the prediction intervals surpasses the threshold of $1-\alpha$ for different significance levels $\alpha\in\{0.05,0.1,0.2\}$. Additionally, we report the width of the resulting intervals in Supplement G.
@@ 5.1 Datasets
Medical dataset: We demonstrate the applicability of our CP method to medical datasets on the MIMIC-III dataset [26]. MIMIC-III contains de-identified health records from patients admitted to critical care units at large hospitals. Our goal is to predict patient outcomes in terms of blood pressure when treated with a different duration of mechanical ventilation. We use 8 confounders from medical practice (e.g., respiratory rate, hematocrit). Overall, we consider 14,719 patients, split into train (60%), validation (10%), calibration (20%), and test (10%) sets. Details are in Supplement F.
@@ 5.2 Results for synthetic datasets
We consider two synthetic datasets with different propensity scores and outcome functions. Dataset 1 uses a step-wise propensity function and a concave outcome function. Dataset 2 is more complex and uses a Gaussian propensity function and oscillating outcome functions. Both datasets contain a single discrete confounder, a continuous treatment, and a continuous outcome. By choosing low-dimensional datasets, we later render it possible to plot the treatment–response curves so that one can inspect the prediction intervals visually. (We later also show that our method scales to high-dimensional settings as part of the real-world dataset.) Details about the data generation are in Supplement F.
@@
We now evaluate the faithfulness of our CP intervals. On each dataset, we analyze the performance of the prediction intervals in the presence of various soft interventions $\Delta_{a}\in\{1,5,10\}$ and hard interventions $a^{\ast}\in\{7x,10x\}$ for each $X=x$. We average the empirical coverage across 50 runs with random seeds. The results are in Fig. 4 and Fig. 5. Additionally, we report the empirical coverage of the GP.
We make the following observations. First, our CP intervals comply with the targeted significance level $\alpha$ and are therefore faithful. Second, both MC dropout and the deep ensemble method have a considerably lower coverage, implying that the intervals are not faithful. This is in line with the literature, where MC dropout is found to produce poor approximations of the posterior [34]. In particular, the ensemble method is highly unfaithful. Thus, we will not consider this baseline in all the following experiments. Third, our method has only a small
@@
variability in terms of empirical coverage, whereas the empirical coverage of MC dropout varies greatly. This corroborates the robustness of our method. Fourth, the results are consistent for both datasets. Fifth, the GP is only able to capture the true potential outcome in the prediction intervals for small distribution shifts ($\Delta=1$) on dataset 2. However, the empirical coverage is extremely low: $\alpha=0.05$: 0.125; $\alpha=0.1$: 0.125; $\alpha=0.2$: 0.0833. This aligns with our expectations, as the aleatoric uncertainty in our experiments is low. Therefore, the GP intervals are small (average width of 0.1293) and barely valid after the intervention. In sum, this demonstrates our CP method’s effectiveness.
@@
Table 1: Coverage of the intervals from our CP method / MC dropout for various hard interventions $a^{\ast}$ and significance levels $\alpha$. Intervals with coverage $\geq 1-\alpha$ are considered faithful (bold numbers).
@@
Table 1 presents the empirical coverage of the intervals from our CP method vs. MC dropout across different $\alpha$ and hard interventions $a^{\ast}$. We observe that our CP intervals are effective and achieve the intended coverage. In contrast, MC dropout does not provide faithful intervals. Our findings are again in line with the literature, where MC dropout is found to produce poor approximations of the posterior and thus might provide poor coverage [34]. We present further results in Supplement G.
Insights: We plot the intervals across different levels $\alpha$ and covariates $X$ (Fig. 6), allowing us to inspect the intervals visually. We observe that the intervals behave as expected: they become sharper with increasing significance level $\alpha$. We further see that our CP intervals are slightly wider (see details in Supplement G), yet this is intended as it ensures that the intervals are faithful. Our CP intervals (blue) generally include the true outcome. In contrast, the intervals from MC dropout (orange) often do not include the true outcome (e.g., see the bottom row Fig. 6) and are thus not faithful.

### v4

@@ 2 Related Work
Several extensions have been developed for CP. One literature stream focuses on CP with marginal coverage under distribution shifts between training and test data [e.g., 8, 11, 15, 16, 17, 18, 19, 36, 42, 49, 56]. Our setting later also involves a distribution shift due to the intervention on the treatment but differs from the latter in that the true distribution shift is unknown. Another literature stream constructs intervals conditional on the variables following the shifted distribution. Since, in general, exact conditional coverage has been proven impossible [35, 51], the works in this literature stream have two key limitations: (1) they only guarantee approximate conditional coverage [e.g., 5, 7, 35, 44]; or (2) they are restricted to specific data structures such as binary variables [e.g., 35, 51]. Because of that, none of the existing methods for marginal and conditional coverage can be applied to derive prediction intervals with finite-sample validity guarantees for causal quantities of continuous treatments.
Conformal prediction for causal quantities: Only a few works focus on CP for causal quantities (see Fig. 2). Examples are methods aimed at off-policy learning [48, 59], conformal sensitivity analysis [57], or meta-learners for the conditional average treatment effect (CATE) [2, 27, 36, 53]. However, there are crucial differences to our setting: First, the existing works (a) assume that the propensity is known and thus achieve finite-sample coverage guarantees, or the existing works (b) focus on the easier task of giving asymptotic guarantees but then might suffer from under-coverage because of which the intervals are not faithful. Only Lei and Candès [36] provides finite-sample coverage guarantees under estimated propensity scores. However, all existing CP methods are designed for binary or discrete treatments. Applying such methods to discretized continuous treatments leads to ill-defined causal estimands. Therefore, none of the existing methods are applicable to our continuous treatment setting. We offer a detailed discussion in Supplement F.
@@ 3 Problem formulation
Second, the propensity score $\pi$ is commonly unknown in observational data and, therefore, must be estimated, which introduces additional uncertainty that one must account for when constructing CP intervals ($\rightarrow$ Challenge  b). Crucially, existing coverage guarantees [e.g., 51, 49] do not hold in our setting. Instead, we must derive new intervals with valid coverage under propensity shift.
@@ 4 CP intervals for potential outcomes of continuous treatments
$\hat{\theta}_{S}:=\operatorname*{arg\,min}\limits_{\theta\in\mathbb{R}}\frac{1}{n-m}\left(\sum_{i=m+1}^{n}l_{\alpha}(\theta,S_{i})+l_{\alpha}(\theta,S)\right),$ (3)
@@
$l_{\alpha}(\theta,S):=(\alpha-\mathbf{1}_{[\theta-S<0]})(\theta-S)=\begin{cases}(1-\alpha)(S-\theta),&\text{ if }S\geq\theta,\\ \alpha(\theta-S),&\text{ if }S<\theta.\end{cases}$ (4)
@@ 4.1 Scenario 1: Known propensity score
Let $\mathcal{F}$ define a finite-dimensional function class that includes the function $f$ characterizing the shift in the (potentially unknown) propensity function $\pi$ (see Eq. 2). Define the distribution-shift-calibrated $(1-\alpha)$-quantile of the non-conformity scores as
$\hat{g}_{S}(X_{n+1}):=\operatorname*{arg\,min}\limits_{g\in\mathcal{F}}\frac{1}{n-m}\left(\sum_{i=m+1}^{n}l_{\alpha}(g(X_{i}),S_{i})+l_{\alpha}(g(X_{n+1}),S)\right)$ (5)
@@
Building upon Lemma 4.1, we derive our first main result in Theorem 4.2. We define the finite-dimensional function class of interest as $\mathcal{F}:=\{\theta\frac{\pi(a+\Delta_{A}\mid x)}{\pi(a\mid x)}\mid\theta\in\mathbb{R}^{+}\}$. It is easy to verify that all $f\in\mathcal{F}$ represent the desired propensity shift to $\tilde{\pi}(a\mid x)=\pi(a+\Delta_{A}\mid x)$ as defined in Eq. 2.
However, note that the optimization problem in Eq. 5 requires knowledge about the true or optimal imputed scores $S_{n+1}$. Directly solving the problem in this form would require running it for all possible imputed values $S\in\mathbb{R}$, i.e., an infinite amount of times. As a remedy, we exploit the strong duality property and present our results in terms of a dual problem formulation.
@@
Consider a new datapoint with $X_{n+1}=x_{n+1}$, $A_{n+1}=a_{n+1}$, and $A^{\ast}(X_{n+1})=a^{*}=a_{n+1}+\Delta_{A}$. Let $\eta^{S}=\{\eta_{m+1}^{S},\ldots,\eta_{n+1}^{S}\}\in\mathbb{R}^{n-m}$ be the optimal solution to
@@
$\text{ s.t.}\qquad-\alpha\leq\eta_{i}\leq 1-\alpha,\quad\forall i=m+1,\ldots,n+1,$
@@
We provide a full proof in Supplement D.2. Here, we briefly outline the underlying idea of the proof. First, we show that the function class $\mathcal{F}$ indeed satisfies Eq. equation 2 for the intervention $A^{*}(X)=A+\Delta_{A}$, and we then rewrite Eq. equation 5 as a convex optimization problem. Next, we exploit the strong duality property. We optimize over the corresponding dual problem to receive a dual prediction set with equal coverage probability. Finally, we derive $S^{\ast}$ from the dual prediction set to construct $C_{n+1}$ and prove the overall coverage guarantee. ∎
@@ 4.2 Scenario 2: Unknown treatment policy
$\hskip-8.5359pt\tilde{\pi}(a\mid x)=\frac{\delta_{a^{\ast}}(a)}{\hat{\pi}(a\mid x)}\frac{\hat{\pi}(a\mid x)}{\pi(a\mid x)}\pi(a\mid x)=c_{a}\frac{\delta_{a^{\ast}}(a)}{\hat{\pi}(a\mid x)}\pi(a\mid x)=\frac{f(a,x)}{\mathbb{E}_{P}[f(A,X)]}\pi({a\mid x}),$ (13)
@@
$\hskip-7.11317pt\mathcal{F}:=\Bigg\{\frac{c_{a}}{\sqrt{2\pi}\sigma}\frac{\exp(-\frac{(a-a^{\ast})^{2}}{2\sigma^{2}})}{\hat{\pi}(a\mid x)}\Bigm|0<\sigma,\frac{1}{M}\leq c_{a}\leq M\Bigg\}.$ (16)
@@
can be represented in terms of Eq. equation 2 through functions $f\in\mathcal{F}$.
@@
for the imputed score $S_{n+1}=S$. As the score is unknown, computing the CP interval would require solving equation 18 for all $S\in\mathbb{R}$, yet which is computationally infeasible. Before, we exploited properties of the dual optimization problem and the Lagrange multipliers of the convex problem in Theorem 4.2 to efficiently compute the CP intervals. However, the present non-convex problem does not automatically allow for the same simplifications. Instead, we now present a remedy for efficient computation of the CP intervals in the following lemma. We prove Lemma 4.4 in Supplement D.
@@
The problem equation 18 is Type-I invex and satisfies the linear independence constraint qualification (LICQ).
Lemma 4.4 allows us to derive properties of the present non-convex optimization problem in terms of the Karush-Kuhn-Tucker (KKT) conditions. For this, we note that the fulfillment of the LICQ serves as a sufficient regularity condition for the KKT to hold at any (local) optimum of equation 18. Combined with the Type-I invexity of the objective function and the constraints, the KKT conditions are not only necessary but also sufficient for a global optimum. As a result, we can employ the KKT conditions at the optimal values [^6] We provide an interpretation of the optimal values in Supplement G. Therein, we further discuss the implications of the proposed kernel smoothing of $\delta_{a^{\ast}}(a)$. $\sigma^{\ast}$ and $c_{a}^{\ast}$ to derive coverage guarantees of our CP interval in a similar fashion as in Theorem 4.2. We thus arrive at the following theorem to provide CP intervals for the scenario with unknown propensity scores.
@@ 5 Experiments
Baselines: As we have discussed above, there are no baselines that directly compute prediction intervals with finite-sample guarantees for potential outcomes of continuous treatments. Therefore, we compare our method against MC dropout [14] and deep ensemble methods [33]. Yet, we again emphasize that MC dropout is an ad hoc method with poor approximations of the posterior, which is known to give unfaithful intervals [34]. Furthermore, we report the empirical coverage achieved by intervals from a Gaussian process regression (GP). By doing so, we consider a method that assesses the underlying aleatoric uncertainty. Additionally, we compare our method against the naive vanilla CP and the method by [36] for binary treatments in Supplements C and G.
Implementation: All methods are implemented with $\phi$ as a multi-layer perceptron (MLP) and an MC dropout regularization of rate 0.1. Crucially, we use the identical MLP for both our CP method and MC dropout. Hence, all performance gains must be attributed to the coverage guarantees of our conformal method. In the MC dropout baseline, the uncertainty intervals are computed via Monte Carlo sampling. In scenario 2, we perform conditional density estimation by conditional normalizing flows [50]. Implementation and training details are in Supplement G.
Performance metrics: We evaluate the methods in terms of whether the prediction intervals are faithful [e.g., 21]. That is, we compute whether the empirical coverage of the prediction intervals surpasses the threshold of $1-\alpha$ for different significance levels $\alpha\in\{0.05,0.1,0.2\}$. Additionally, we report the width of the resulting intervals in Supplement H.
@@ 5.1 Datasets
Medical dataset: We demonstrate the applicability of our CP method to medical datasets on the MIMIC-III dataset [26]. MIMIC-III contains de-identified health records from patients admitted to critical care units at large hospitals. Our goal is to predict patient outcomes in terms of blood pressure when treated with a different duration of mechanical ventilation. We use 8 confounders from medical practice (e.g., respiratory rate, hematocrit). Overall, we consider 14,719 patients, split into train (60%), validation (10%), calibration (20%), and test (10%) sets. Details are in Supplement G.
@@ 5.2 Results for synthetic datasets
We consider two synthetic datasets with different propensity scores and outcome functions. Dataset 1 uses a step-wise propensity function and a concave outcome function. Dataset 2 is more complex and uses a Gaussian propensity function and oscillating outcome functions. Both datasets contain a single discrete confounder, a continuous treatment, and a continuous outcome.
@@
By choosing low-dimensional data-sets, we later render it possible to plot the treatment–response curves so that one can inspect the prediction intervals visually. (We later also show that our method scales to high-dimensional settings as part of the real-world dataset.) Details about the data generation are in Supplement G.
We evaluate the faithfulness of our CP intervals. On each dataset, we analyze the performance of the intervals in the presence of various soft interventions $\Delta_{a}\in\{1,5,10\}$ and hard interventions $a^{\ast}\in\{7x,10x\}$ for each $X=x$. We average the empirical coverage across 50 runs with random seeds. The results are in Fig. 4 and Fig. 5. Additionally, we report the empirical coverage of the GP.
@@
Table 1: Coverage of the intervals from our CP method / MC dropout for various hard interventions $a^{\ast}$ and significance levels $\alpha$. Intervals with coverage $\geq 1-\alpha$ are considered faithful.
We make the following observations. First, our CP intervals comply with the targeted significance level $\alpha$ and are therefore faithful. Second, both MC dropout and the deep ensemble method have a considerably lower coverage, implying that the intervals are not faithful. This is in line with the literature, where MC dropout is found to produce poor approximations of the posterior [34]. In particular, the ensemble method is highly unfaithful. Thus, we will not consider this baseline in all the following experiments. Third, our method has only a small variability in terms of empirical coverage, whereas the empirical coverage of MC dropout varies greatly. This corroborates the robustness of our method. Fourth, the results are consistent for both datasets. Fifth, the GP is only able to capture the true potential outcome in the prediction intervals for small distribution shifts ($\Delta=1$)
@@
on dataset 2. However, the empirical coverage is extremely low: $\alpha=0.05$: 0.125; $\alpha=0.1$: 0.125; $\alpha=0.2$: 0.0833. This aligns with our expectations, as the aleatoric uncertainty in our experiments is low. Therefore, the GP intervals are small (average width of 0.1293) and barely valid after the intervention. In sum, this demonstrates our CP method’s effectiveness.
Table 1 presents the empirical coverage of the intervals from our CP method vs. MC dropout across different $\alpha$ and hard interventions $a^{\ast}$. We observe that our CP intervals are effective and achieve the intended coverage. In contrast, MC dropout does not provide faithful intervals. Our findings are again in line with the literature, where MC dropout is found to produce poor approximations of the posterior and thus might provide poor coverage [34]. We present further results in Supplement H.
Insights: We plot the intervals across different levels $\alpha$ and covariates $X$ (Fig. 6), allowing us to inspect the intervals visually. We observe that the intervals behave as expected: they become sharper with increasing significance level $\alpha$. We further see that our CP intervals are slightly wider (see details in Supplement H), yet this is intended as it ensures that the intervals are faithful. Our CP intervals (blue) generally include the true outcome. In contrast, the intervals from MC dropout (orange) often do not include the true outcome (e.g., see the bottom row Fig. 6) and are thus not faithful.